A company uses the periodic inventory system and had the following activity during the current monthly period.

November 1: Beginning inventory 108 Units @ $15
November 5: Purchased 108 Units @ $30
November 8: Purchased 58 Units @ $30
November 16: Sold 158 Units @ $85
November 19: Purchased 55 Units @ $30


Using the weighted-average inventory method, the company's ending inventory would be:

  • $3,270

  • $3,390

  • $5,010

  • $4,288

  • $1,620
Answer #1
Computation of ending inventory using weighted average
Nov-01 Beginning inventory 108 15 1620
Nov-05 purchases 108 30 3240
Nov-08 purchased 58 30 1740
Nov-19 purchased 55 30 1650
Total 329 8250
Cost per unit = total cost/total unit
Cost per unit = 8250/329 = 25.08
Inventory at the end = (329-158)*25.08
Inventory at the end = 4288
Optipon " $4288 is CORRECT